# Counting Equilibria of the Electrostatic Potential

### v1

@@ C.2. Rotated prisms
C.3. The square anti-prism
We shall now conduct a more thorough study of the square anti-prism. For concreteness, let $A_{1}=(1,0,\frac{h}{2})$, $A_{2}=(0,1,\frac{h}{2})$, $A_{3}=(-1,0,\frac{h}{2})$, $A_{4}=(0,-1,\frac{h}{2})$ be the vertices of the square at the top, and $B_{1}=(\frac{\sqrt{2}}{2},\frac{\sqrt{2}}{2},-\frac{h}{2})$, $B_{2}=(-\frac{\sqrt{2}}{2},\frac{\sqrt{2}}{2},-\frac{h}{2})$, $B_{3}=(-\frac{\sqrt{2}}{2},-\frac{\sqrt{2}}{2},-\frac{h}{2})$, $B_{4}=(\frac{\sqrt{2}}{2},-\frac{\sqrt{2}}{2},-\frac{h}{2})$ the vertices of the square at the bottom. Hence, the origin of $\mathbb{R}^{3}$ is the center, the $x_{3}$-axis is the axis, $R=1$ is the radius, and $h$ is the height of the anti-prism. Its lateral edges are the ones that connect vertices of the two squares.
Proposition 4.
For $h<\sqrt{2}R$, there is an equilibrium on the segment connecting the center of the square anti-prism to the midpoint of every lateral edge.
@@ C.3. The square anti-prism
Consider the open half-lines that emanate from the origin and pass through the midpoints of the lateral edges of the anti-prism. There is no loss of generality in only considering one of these, namely the one that maps $t>0$ to $L(t)=\frac{t}{2}(A_{1}+B_{1})=\frac{t}{4}(2+\sqrt{2},\sqrt{2},0)$, with normal vectors $n_{1}=(0,0,1)$ and $n_{2}=(-\sqrt{2},2+\sqrt{2},0)$. Then, by direct computation, we find that along $L$ the directional derivatives both vanish:
(44) $\langle n_{1},\nabla V(L(t))\rangle$ $=\frac{\partial V}{\partial x_{3}}(L(t))=0,$
(45) $\langle n_{2},\nabla V(L(t))\rangle$ $=-\sqrt{2}\frac{\partial V}{\partial x_{1}}(L(t))+(2+\sqrt{2})\frac{\partial V}{\partial x_{2}}(L(t))=0.$
Therefore, along $L$, the gradient $\nabla V$ is parallel to it, and
(46) $\langle\dot{L}(t),\nabla V(L(t))\rangle$ $=\frac{2+\sqrt{2}}{4}\frac{\partial V}{\partial x_{1}}(L(t))+\frac{\sqrt{2}}{4}\frac{\partial V}{\partial x_{2}}(L(t)),$
which we can also compute explicitly but has a very long and not very enlightening expression. Instead, notice that this quantity vanishes when $t=0$ and its first derivative at this point is
(47) $\frac{d}{dt}\Big{|}_{t=0}\langle\dot{L}(t),\nabla V(L(t))\rangle=16\frac{2\sqrt{2}+5}{(h^{2}+4)^{5/2}}(2-h^{2}),$
which is positive if $h<\sqrt{2}R$. In particular, there is $\varepsilon>0$ such that this function is positive for $t\in(0,\varepsilon)$ and negative for $t\in(-\varepsilon,0)$. On the other hand, evaluating this function at $t=\pm 1$, we find $\langle\dot{L}(-1),\nabla V(L(-1))\rangle>0$ and $\langle\dot{L}(1),\nabla V(L(1))\rangle<0$. By the intermediate value theorem, there must be two extra zeroes along this line. ∎
@@
C.4. The triangular anti-prism
We now turn to the case of the triangular anti-prism, using the geometric realization in which the vertices at the top are $A_{1}=\left(\sqrt{3},1,\frac{h}{2}\right)$, $A_{2}=\left(-\sqrt{3},1,\frac{h}{2}\right)$, $A_{3}=\left(0,-2,\frac{h}{2}\right)$, and those at the bottom are $B_{1}=\left(\sqrt{3},-1,-\frac{h}{2}\right)$, $B_{2}=\left(0,2,-\frac{h}{2}\right)$, $B_{3}=\left(-\sqrt{3},-1,-\frac{h}{2}\right)$. Hence, $0\in\mathbb{R}^{3}$ is the center, the $x_{3}$-axis is the axis, $R=2$ is the radius, and $h$ is the height of the anti-prism. We start with a result analogous to the proposition we proved for the square anti-prism.
Proposition 5.
For $h<\sqrt{2}R$, there is an equilibrium on the open half-line connecting the center of the triangular anti-prism to the midpoint of every lateral edge.
Proof.
Without loss of generality, we can consider the lateral edge connecting $A_{1}$ and $B_{1}$. Its midpoint is $\frac{1}{2}(A_{1}+B_{1})=(\sqrt{3},0,0)$, so the half-line connecting the origin to this point is the positive half of the $x_{1}$-axis. We parametrize it by $L(t)=t(\sqrt{3},0,0)$ and note that is has normal vectors $n_{1}=(0,1,0)$ and $n_{2}=(0,0,1)$. Computing the derivatives in the direction of $n_{1}$ and $n_{2}$ along this line, we find that both vanish:
(49) $\langle\nabla V(L(t)),n_{1}\rangle$ $=\frac{\partial V}{\partial x_{2}}(L(t))=0\mbox{\rm~{}~{}and~{}~{}}\langle\nabla V(L(t)),n_{2}\rangle=\frac{\partial V}{\partial x_{3}}(L(t))=0.$
The remaining directional derivative along $\dot{L}(t)=(\sqrt{3},0,0)$ is
(50) $\langle\dot{L}(t),\nabla V(L(t))\rangle$ $=\sqrt{3}\frac{\partial V}{\partial x_{1}}(L(t))$
(51) $=-2\sqrt{3}\frac{t+1}{\left(3(t+1)^{2}+1+\frac{h^{2}}{4}\right)^{3/2}}-2\sqrt{3}\frac{t-1}{\left(3(t-1)^{2}+1+\frac{h^{2}}{4}\right)^{3/2}}$
(52) $\ \ \ \ -16\sqrt{3}\frac{t}{\left(12t^{2}+16+h^{2}\right)^{3/2}},$
and the derivative of this quantity at $t=0$ is given by $48\sqrt{3}\frac{8-h^{2}}{(h^{2}+16)^{5/2}}$, which is positive if $h<2\sqrt{2}=\sqrt{2}R$ and so it is positive for small positive $t$. On the other hand, we have
(53) $\sqrt{3}\frac{\partial V}{\partial x_{1}}(L(1))=-\frac{4\sqrt{3}}{\left(13+\frac{h^{2}}{4}\right)^{3/2}}-\frac{16\sqrt{3}}{\left(h^{2}+28\right)^{3/2}}<0,$
so it must have a zero in the interval $(0,1)$ corresponding to an equilibrium. ∎
Again we compute the Taylor expansion at the origin as a way to understand the degenerate critical point there for $h=\sqrt{2}R$.
@@ C.4. The triangular anti-prism
Appendix D Proofs of Propositions 1 and 2

### v2

@@ C.2. Rotated prisms
C.3. Existence results for equilibria in ati-prisms
We shall consider an anti-prims whose top and bottom faces are $N$-gons and denote their minimum angle of symmetry by $\alpha=\frac{2\pi}{N}$. We shall let $R>0$ be the radius of these $N$-gons and $h>0$ the height of the anti-prism, then its relative height is defined by $\frac{h}{R}$.
Identify $\mathbb{R}^{3}\cong\mathbb{C}\times\mathbb{R}$, then for $\ell\in\{1,\ldots,N\}$ we shall denote the vertices of the top $N$-gon as $A_{\ell}=(Re^{i(\ell\alpha+\frac{\alpha}{4})},\frac{h}{2})$ and those at the bottom $N$-gon as $B_{\ell}=(Re^{-i(\ell\alpha+\frac{\alpha}{4})},-\frac{h}{2})$. Then, the electric potential generated by placing unit point charges at the vertices of this anti-prism is
$V(x)=V_{u}(x)+V_{d}(x),$
where
$V_{u}(x)=\sum_{\ell=1}^{N}\frac{1}{|x-A_{\ell}|},\ \ V_{d}(x)=\sum_{\ell=1}^{N}\frac{1}{|x-B_{\ell}|}.$
We shall now state and prove the following result.
Theorem 5.
Let $V$ be the electric potential generated by placing unit point charges at the vertices of an anti-prism as above and suppose its relative height satisfies $\frac{h}{R}<\sqrt{2}$. Then, there are electrostatic points located along the line segments connecting the center of the anti-prism to the midpoints of its lateral edges.
@@ C.3. Existence results for equilibria in ati-prisms
We shall consider the case of the line segment connecting the center to the midpoint of the edge connecting $A_{N}$ to $B_{N}$. The remaining cases can be handled similarly.
The midpoint between $A_{N}$ and $B_{N}$ is $\frac{A_{N}+B_{N}}{2}=(R\cos(\frac{\alpha}{4}),0,0)$ and we can parametrize the segment that connects the origin to it using $t\in[0,1]\mapsto(tR\cos(\frac{\alpha}{4}),0,0)$. The derivatives of $V(x)$ in the normal directions $n_{1}=(0,1,0)$ and $n_{2}=(0,0,1)$ are given by
$\langle\nabla V,n_{1}\rangle$ $=\frac{\partial V_{u}}{\partial x_{2}}+\frac{\partial V_{d}}{\partial x_{2}}$
$\langle\nabla V,n_{2}\rangle$ $=\frac{\partial V_{u}}{\partial x_{3}}+\frac{\partial V_{d}}{\partial x_{3}}.$
Now, we compute each of the terms in the rights hand side in turn. We start with
$\frac{\partial V_{u}}{\partial x_{2}}$ $=-\sum_{\ell=1}^{N}\frac{(x_{2}-R\sin(\ell\alpha+\frac{\alpha}{4}))}{|x-A_{\ell}|^{3}}$
$\frac{\partial V_{d}}{\partial x_{2}}$ $=-\sum_{\ell=1}^{N}\frac{(x_{2}+R\sin(\ell\alpha+\frac{\alpha}{4}))}{|x-B_{\ell}|^{3}},$
on the other hand, along the line connecting 00 to $\frac{A_{N}+B_{N}}{2}$, we have $x_{2}=0=x_{3}$ and
$|x-A_{\ell}|$ $=|(x_{1}-R\cos(\ell\alpha+\tfrac{\alpha}{4}),R\sin(\ell\alpha+\tfrac{\alpha}{4}),0)|$
$=|(x_{1}-R\cos(-\ell\alpha-\tfrac{\alpha}{4}),R\sin(-\ell\alpha-\tfrac{\alpha}{4}),0)|$
$=|x-B_{\ell}|.$
which shows that along such a line
$\frac{\partial V_{u}}{\partial x_{2}}=-\frac{\partial V_{d}}{\partial x_{2}}.$
On the other hand,
$\frac{\partial V_{u}}{\partial x_{3}}$ $=-\sum_{\ell=1}^{N}\frac{(x_{3}-\frac{h}{2})}{|x-A_{\ell}|^{3}}$
$\frac{\partial V_{d}}{\partial x_{3}}$ $=-\sum_{\ell=1}^{N}\frac{(x_{3}+\frac{h}{2})}{|x-B_{\ell}|^{3}},$
and, again, along the line $x=2=0=x_{3}$ we have
$\frac{\partial V_{u}}{\partial x_{3}}=-\frac{\partial V_{d}}{\partial x_{3}}.$
Hence,
$\langle\nabla V,n_{1}\rangle=0=\langle\nabla V,n_{2}\rangle.$
Now, we shall finally consider the directional derivative in the direction of the tangent vector to the line under consideration $v=(1,0,0)$
$\langle\nabla V,v\rangle$ $=\frac{\partial V}{\partial x_{1}}$
$=\frac{\partial V_{u}}{\partial x_{1}}+\frac{\partial V_{d}}{\partial x_{1}}$
$=-\sum_{\ell=1}^{N}\left(\frac{(x_{1}-R\cos(\ell\alpha+\frac{\alpha}{4}))}{|x-A_{\ell}|^{3}}+\frac{(x_{1}-R\cos(-\ell\alpha-\frac{\alpha}{4}))}{|x-B_{\ell}|^{3}}\right).$
Furthermore, along the line $x_{2}=0=x_{3}$, we have
$\langle\nabla V,v\rangle$ $=-\sum_{\ell=1}^{N}\left(\frac{(x_{1}-R\cos(\ell\alpha+\frac{\alpha}{4}))}{|x-A_{\ell}|^{3}}+\frac{(x_{1}-R\cos(-\ell\alpha-\frac{\alpha}{4}))}{|x-B_{\ell}|^{3}}\right)$
$=-2\sum_{\ell=1}^{N}\frac{(x_{1}-R\cos(\ell\alpha+\frac{\alpha}{4}))}{|x-A_{\ell}|^{3}}$
and inserting $x_{1}=tR\cos(\frac{\alpha}{4})$ gives
$\langle\nabla V,v\rangle$ $=-\frac{2}{R^{2}}\sum_{\ell=1}^{N}\frac{t\cos(\frac{\alpha}{4})-\cos(\ell\alpha+\frac{\alpha}{4})}{\left((t\cos(\frac{\alpha}{4})-\cos(\ell\alpha+\frac{\alpha}{4}))^{2}+(\sin(\ell\alpha+\frac{\alpha}{4}))^{2}+\frac{h^{2}}{4}\right)^{\frac{3}{2}}}.$
Now, notice that $\cos(\frac{\alpha}{4})\geq\cos(\ell\alpha+\frac{\alpha}{4})$ with strict inequality for $\ell\neq N$, which shows that for $t=1$ we have $\langle\nabla V,v\rangle<0$. On the other hand, for evaluating this quantity near $t=0$, we compute its Taylor series at $t=0$
$\langle\nabla V,v\rangle$ $=\frac{16}{R^{2}}\frac{1}{\left((4R^{2}+h^{2}\right)^{\frac{3}{2}}}\sum_{\ell=1}^{N}\cos(\ell\alpha+\frac{\alpha}{4})$
$\ \ \ \ +\frac{16t}{R^{2}}\frac{\cos(\frac{\alpha}{4})}{(4R^{2}+h^{2})^{\frac{5}{2}}}\sum_{\ell=1}^{N}\left(12R^{2}\cos(\ell\alpha+\frac{\alpha}{4})^{2}-4R^{2}-h^{2}\right)+O(t^{2}).$
Now, using the facts that
$\sum_{\ell=1}^{N}\cos(\ell\alpha+\frac{\alpha}{4})=0,\ \ \sum_{\ell=1}^{N}\cos(\ell\alpha+\frac{\alpha}{4})^{2}=\frac{N}{2},$
we find that
$\langle\nabla V,v\rangle$ $=\frac{16t}{R^{2}}\frac{\cos(\frac{\alpha}{4})}{(4R^{2}+h^{2})^{\frac{5}{2}}}\left(6NR^{2}-4NR^{2}-Nh^{2}\right)+O(t^{2})$
$=\frac{16Nt}{R^{2}}\frac{\cos(\frac{\alpha}{4})}{(4R^{2}+h^{2})^{\frac{5}{2}}}\left(2R^{2}-h^{2}\right)+O(t^{2}).$
Hence, if $h<\sqrt{2}h$ we have that along this line $\langle\nabla V,v\rangle$ is positive for small positive $t>0$. Combining this with the previously proven fact that it is negative for $t=1$ we find, from the intermediate value theorem, that $\langle\nabla V,v\rangle$ has a zero at some $t\in(0,1)$. This finishes the proof that $\nabla V$ as a zero along the segment connecting 00 to $\frac{A_{N}+B_{N}}{2}$. ∎
@@
Again, we compute the Taylor expansion at the origin as a way to understand the degenerate critical point there for $h=\sqrt{2}R$.
@@
Appendix D Proofs of Propositions 2 and 3